### v2

@@ 6. Conformal Prediction Regions as Images of the Imprecise Highest Density Region Functor
[…]
@@
Because $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$, the same inequality holds with $\mathcal{M}_{2}$ replaced by $\mathcal{M}_{1}$, so $A$ belongs to the family of sets whose intersection defines $B_{\mathcal{M}_{1}}^{\alpha}$. Consequently $y\in A$, contradicting $y\notin A$. Therefore $y\in B_{\mathcal{M}_{2}}^{\alpha}$, and $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}.$ Hence $\mathsf{IHDR}_{\alpha}$ maps the arrow $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ to the arrow $B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha}$.
@@
Replacing $B_{\mathcal{M}}^{\alpha}$ by its complement $Y\setminus B_{\mathcal{M}}^{\alpha}$ yields a contravariant functor. Allowing $\alpha$ itself to vary introduces a functor $\mathsf{IHDR}\colon[0,1]\times\mathscr{C}\rightarrow\Sigma_{Y}$ that is antitone in the first argument and covariant in the second.
@@
Viewing the conformal-prediction map $\mathsf{IHDR}_{\alpha}$ as a covariant functor immediately yields practical AI privacy benefits. Suppose that we have many sources producing credal sets $\mathcal{M}_{j}$. First, any differentially-private or otherwise noise-adding transformation applied to each site’s credal set $\mathcal{M}_{j}$ propagates through the functor without violating the coverage guarantee, so statistical validity is preserved “for free”.
@@ 7. Further Properties of our Categories
In this section, we ask ourselves whether we can further characterize the categories that we introduced earlier in Sections 3 and 4. Throughout the section, if restrictions have to be imposed on the objects or morphisms of a category, we will report them as superscript and subscripts of the ensuing subcategory, respectively.
@@ 7.1. Monoidal Categories
are homeomorphisms. They define morphisms in $\mathbf{UHCont}_{\text{c}}$ via single-valued, continuous correspondences (e.g., $(x,*)\mapsto x$) with compact singleton values.
@@
(4) Coherence Conditions. The coherence diagrams (pentagon and triangle) commute strictly, as they correspond to natural identifications of Cartesian products in the category of topological spaces. Thus, the associators and unitors satisfy the coherence conditions. ∎
@@
As perhaps expected from the proof of Theorem 16, the category $\mathbf{WMeas}_{\text{uc}}$ is monoidal.
@@
1. Tensor is a bifunctor. For morphisms $F:X\!\Rightarrow\!Y$ and $G:X^{\prime}\!\Rightarrow\!Y^{\prime}$ the map $(F\otimes G)(x,x^{\prime})=F(x)\times G(x^{\prime})$ is:
@@
Weakly measurable. Let $O=\bigcup_{k\in\mathbb{N}}(B_{k}\times C_{k})$ be an open subset of $Y\times Y^{\prime}$ written as a countable union of basic rectangles (recall that $Y,Y^{\prime}$ are second-countable). Then,
@@
2. Associativity, unit, symmetry. Because $\otimes$ is the cartesian product on objects and the point-wise cartesian product on fibres, we have the canonical homeomorphisms
@@
Each of these maps is single-valued and continuous, hence a morphism in $\mathbf{WMeas}_{\text{uc}}$ (singleton fibres are compact). They play the role of associator, left- and right-unitors, and symmetry, respectively.
3. Coherence. The associator, unitors, and symmetry just introduced are the usual homeomorphisms that witness the cartesian product as a symmetric monoidal structure in Set. Consequently the pentagon, triangle, and hexagon coherence diagrams commute exactly as they do in Set, and the same proofs apply verbatim in $\mathbf{WMeas}_{\text{uc}}$.
@@ 7.2. Functors Between Categories
Functors do exist if we focus on restrictions of $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$ [^12] We defer the study of monoidal functors to future research. In particular, call $\mathbf{WMeas}_{\text{uc,uhc}}\subset\mathbf{WMeas}_{\text{uc}}$ the category whose objects are measurable Polish spaces with the Borel $\sigma$-algebra, and whose morphisms are uniformly compact-valued, weakly measurable, and upper hemicontinuous correspondences. Then, the following holds.
@@
(2) Composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be morphisms in $\mathbf{WMeas}_{\text{uc,uhc}}$. Then,
  • Both are uniformly compact-valued and upper hemicontinuous;
  • Their composition $\Psi\circ\Phi$ is also uniformly compact-valued (Aliprantis and Border,, 2006, Page 558 and (a consequence of) Example 17.24) and upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23);
  • Weak measurability is preserved under composition (by the Proof of Theorem 7).
So $\Psi\circ\Phi$ is a morphism in $\mathbf{WMeas}_{\text{uc,uhc}}$, and $\mathcal{F}(\Psi\circ\Phi)=\mathcal{F}(\Psi)\circ\mathcal{F}(\Phi)$.
@@
Let now $\mathbf{UHCont}_{\text{c}}^{\text{Pa}}\subset\mathbf{UHCont}$ be the category whose objects are Polish topological spaces, and whose morphisms are upper hemicontinuous, uniformly compact-valued correspondences having analytic graphs. Then, the following holds.
Theorem 19 (Faithful Functor between $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}$ and $\mathbf{WMeas}_{\text{uc}}$).
@@
$\mathcal{G}:\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}\rightarrow\mathbf{WMeas}_{\text{uc}}$
@@
First, notice that if $X$ is a Polish space with topology $\tau_{X}$, then $(X,\tau_{X},\mathcal{B}(\tau_{X}))$ is a measurable Polish space, a valid object in $\mathbf{WMeas}_{\text{uc}}$.
Also, let $\Phi:X\rightrightarrows Y$ be a morphism in $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}$, i.e. uniformly compact-valued, upper hemicontinuous, and with analytic graph. Then, since $\Phi$ is compact-valued, it follows that $\Phi$ is weakly measurable (Aliprantis and Border,, 2006, Theorem 18.10). Thus $\Phi$ is a valid morphism in $\mathbf{WMeas}_{\text{uc}}$.
We now verify the functorial properties.
(1) Identity. The identity correspondence $x\mapsto\{x\}$ has analytic graph, is trivially upper hemicontinuous and (uniformly) compact-valued, hence maps to the identity in $\mathbf{WMeas}_{\text{uc}}$.
(2) Composition. The composition of two (uniformly) compact-valued, upper hemicontinuous, analytic-graph correspondences again satisfies all three properties (using standard results from multivalued analysis), so $\mathcal{G}(\Psi\circ\Phi)=\mathcal{G}(\Psi)\circ\mathcal{G}(\Phi)$. [^13] The graph of the composite $G_{\Psi\circ\Phi}=\bigl{\{}(x,z):\exists y,\,(x,y)\in G_{\Phi},\;(y,z)\in G_{\Psi}\bigr{\}}$ is the analytic image (Souslin projection) of the analytic set $G_{\Phi}\times Z\cap X\times G_{\Psi}$; Hence it is analytic.
(3) Faithfulness. Since $\mathcal{G}$ is the identity on morphisms, it is injective on $\operatorname{Hom}$-sets. ∎
Notice that the analytic graph condition in Theorem 19 is essential. Upper hemicontinuity alone does not imply weak measurability; Hence, $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}$ is the minimal natural subcategory for which functor $\mathcal{G}$ is defined.
@@ 7.3. Monads
Theorem 21 (Vietoris Monad on $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$).
There exists a monad $(T,\eta_{X},\nu_{X})$ on the category $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$, where
  • $T(X)\coloneqq\mathcal{K}(X)$, the space of compact subsets of $X$, endowed with the Vietoris topology;
  • For a morphism $\Phi:X\rightrightarrows Y$, define $T(\Phi):\mathcal{K}(X)\rightrightarrows\mathcal{K}(Y)$ by
    $T(\Phi)(K)\;\coloneqq\;\bigl{\{}\,L\in\mathcal{K}(Y)\;:\;L\subset\Phi[K]\bigr{\}},\quad\Phi[K]\coloneqq\bigcup_{x\in K}\Phi(x);$
  • The unit $\eta_{X}:X\rightarrow\mathcal{K}(X)$ is given by $\eta_{X}(x)=\{x\}$;
  • The multiplication $\nu_{X}:\mathcal{K}(\mathcal{K}(X))\rightarrow\mathcal{K}(X)$ is defined by
    $\nu_{X}(\mathcal{A})\coloneqq\bigcup_{A\in\mathcal{A}}A,$
    where $\mathcal{K}(\mathcal{K}(X))$ is the hyperspace of nonempty, compact collections of compact subsets of $X$.
@@
We verify the monad axioms.
(1) Well-definedness.
  • $T(X)$ is a compact Hausdorff space (classical result).
  • We now verify that $T(\Phi)$ is a morphism in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. By Lemma 3, $\Phi$ has a closed graph; It also has compact values. It follows from standard results on the Vietoris topology (Michael,, 1951) that
    • The map $K\mapsto T(\Phi)(K)$ is upper hemicontinuous;
    • The values $T(\Phi)(K)$ are nonempty, compact subsets of $\mathcal{K}(Y)$, hence compact in the Vietoris topology.
    Thus, $T(\Phi)$ is again a compact-valued, upper hemicontinuous correspondence. Finally, for morphisms $\Phi:X\!\rightrightarrows\!Y$ and $\Psi:Y\!\rightrightarrows\!Z$ we can see that

### v3

@@ 5. Conformal Prediction Regions as Images of the Imprecise Highest Density Region Functor
$\mathscr{C}\coloneqq\bigl{\{}\mathcal{M}\subseteq\Delta_{Y}:\mathcal{M}\text{ is convex, nonempty, and weak${}^{\star}$-closed}\bigr{\}},$
@@
Because $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$, the same inequality holds with $\mathcal{M}_{2}$ replaced by $\mathcal{M}_{1}$, so $A$ belongs to the family of sets whose intersection defines $B_{\mathcal{M}_{1}}^{\alpha}$. Consequently $y\in A$, contradicting $y\notin A$. Therefore $y\in B_{\mathcal{M}_{2}}^{\alpha}$, and $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}.$
Since $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$ implies $\overline{P}_{\mathcal{M}_{1}}(A)\leq\overline{P}_{\mathcal{M}_{2}}(A)$ for all $A\in\Sigma_{Y}$, we have $\underline{P}_{\mathcal{M}_{1}}(A)=1-\overline{P}_{\mathcal{M}_{1}}(A^{c})\geq 1-\overline{P}_{\mathcal{M}_{2}}(A^{c})=\underline{P}_{\mathcal{M}_{2}}(A)$. Hence
$\{A\in\Sigma_{Y}:\underline{P}_{\mathcal{M}_{2}}(A)\geq 1-\alpha\}\subseteq\{A\in\Sigma_{Y}:\underline{P}_{\mathcal{M}_{1}}(A)\geq 1-\alpha\},$
and therefore $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}$.
Intuitively, enlarging a credal set can only decrease the lower probability (by the monotonicity of lower probabilities), so fewer events clear the $1-\alpha$ bar; Intersecting fewer sets yields a larger IHDR. Hence $\mathsf{IHDR}_{\alpha}$ maps the arrow $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ to the arrow $B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha}$.
@@
Replacing $B_{\mathcal{M}}^{\alpha}$ by its complement $Y\setminus B_{\mathcal{M}}^{\alpha}$ yields a contravariant functor. Allowing $\alpha$ itself to vary introduces a functor $\mathsf{IHDR}\colon[0,1]\times\mathscr{C}\rightarrow\Sigma_{Y}$ that is antitone in the first argument and covariant in the second. $\triangle$
@@
Viewing the conformal prediction map $\mathsf{IHDR}_{\alpha}$ as a covariant functor immediately yields practical AI privacy benefits. Suppose that we have many sources producing credal sets $\mathcal{M}_{j}$. First, differentially-private or otherwise noise-adding transformation applied to each site’s credal set $\mathcal{M}_{j}$ propagates through the functor without violating the coverage guarantee, so statistical validity is preserved “for free”.
@@
Caveat. The functorial guarantee is monotone in the credal set: To preserve (and possibly enlarge) coverage after privatization, each site should output a superset $\widetilde{\mathcal{M}}_{j}\supseteq\mathcal{M}_{j}$ (e.g., add a weak buffer and take closed convex hull). Inner approximations $\widetilde{\mathcal{M}}_{j}\subset\mathcal{M}_{j}$ can shrink $B^{\alpha}_{\mathcal{M}}$ and reduce coverage. For federation, aggregation should be performed via the closed convex hull $\overline{\operatorname{co}}(\cdot)$,
$\mathcal{M}_{\mathrm{agg}}=\overline{\operatorname{co}}\Bigl{(}\bigcup_{j}\widetilde{\mathcal{M}}_{j}\Bigr{)},$
which ensures $B^{\alpha}_{\widetilde{\mathcal{M}}_{j}}\subseteq B^{\alpha}_{\mathcal{M}_{\mathrm{agg}}}$, and thus global validity whenever the true law lies in some $\widetilde{\mathcal{M}}_{j}$.
@@ 6. Further Properties of our Categories
In this section, we ask ourselves whether we can further characterize the categories that we introduced earlier in Sections 3. Throughout the section, if restrictions have to be imposed on the objects or morphisms of a category, we will report them as superscript and subscripts of the ensuing subcategory, respectively.
@@ 6.1. Monoidal Categories
are homeomorphisms. They define morphisms in $\mathbf{UHCont}_{\text{c}}$ via single-valued, continuous correspondences (e.g., $(*,x)\mapsto x$) with compact singleton values.
@@
(4) Bifunctoriality of $\otimes$. For morphisms $\Phi_{1}:X\rightrightarrows X^{\prime}$, $\Psi_{1}:X^{\prime}\rightrightarrows X^{\prime\prime}$ and $\Phi_{2}:Y\rightrightarrows Y^{\prime}$, $\Psi_{2}:Y^{\prime}\rightrightarrows Y^{\prime\prime}$, for each $(x,y)\in X\times Y$,
$\bigl{(}(\Psi_{1}\circ\Phi_{1})\otimes(\Psi_{2}\circ\Phi_{2})\bigr{)}(x,y)=\Bigl{(}\bigcup_{u\in\Phi_{1}(x)}\Psi_{1}(u)\Bigr{)}\times\Bigl{(}\bigcup_{v\in\Phi_{2}(y)}\Psi_{2}(v)\Bigr{)}=\bigcup_{(u,v)\in\Phi_{1}(x)\times\Phi_{2}(y)}\Psi_{1}(u)\times\Psi_{2}(v),$
which equals $\bigl{(}(\Psi_{1}\otimes\Psi_{2})\circ(\Phi_{1}\otimes\Phi_{2})\bigr{)}(x,y)$ by the definition of composition for correspondences. Identities are preserved since $(\mathrm{id}_{X}\otimes\mathrm{id}_{Y})(x,y)=\{x\}\times\{y\}$.
(5) Naturality of the associator and unitors. For morphisms $\Phi_{1}:X\rightrightarrows X^{\prime}$, $\Phi_{2}:Y\rightrightarrows Y^{\prime}$, $\Phi_{3}:Z\rightrightarrows Z^{\prime}$, we have
$\alpha_{X^{\prime},Y^{\prime},Z^{\prime}}\circ\bigl{(}(\Phi_{1}\otimes\Phi_{2})\otimes\Phi_{3}\bigr{)}=\bigl{(}\Phi_{1}\otimes(\Phi_{2}\otimes\Phi_{3})\bigr{)}\circ\alpha_{X,Y,Z},$
since both sides send $(x,y,z)$ to $\Phi_{1}(x)\times\bigl{(}\Phi_{2}(y)\times\Phi_{3}(z)\bigr{)}$ (after re-bracketing). For unitors, naturality reads
$\lambda_{X^{\prime}}\circ(\mathrm{id}_{I}\otimes\Phi_{1})=\Phi_{1}\circ\lambda_{X},\qquad\rho_{Y^{\prime}}\circ(\Phi_{2}\otimes\mathrm{id}_{I})=\Phi_{2}\circ\rho_{Y},$
because $\{*\}\times\Phi_{1}(x)=\Phi_{1}(x)$ and $\Phi_{2}(y)\times\{*\}=\Phi_{2}(y)$.
(6) Coherence Conditions. The pentagon and triangle commute, as in Top with the Cartesian product; Here $\alpha,\lambda,\rho$ are homeomorphisms (single-valued morphisms), so the standard coherence diagrams commute strictly.
(7) Symmetry. The flip map $\sigma_{X,Y}:X\times Y\to Y\times X$, $\sigma(x,y)=(y,x)$, is a homeomorphism, hence a morphism; The usual hexagon coherence holds. Thus $\mathbf{UHCont}_{\mathrm{c}}$ is symmetric monoidal. Moreover, the braiding is natural: $(\Phi_{2}\otimes\Phi_{1})\circ\sigma_{X,Y}=\sigma_{X^{\prime},Y^{\prime}}\circ(\Phi_{1}\otimes\Phi_{2})$. ∎
@@
Recall from Definition 6 that structure $\mathbf{WMeas}_{\text{uc}}$ has compact Polish spaces with the Borel $\sigma$-algebra for objects, and weakly measurable, uniformly compact-valued correspondences for morphisms. $\mathbf{WMeas}_{\text{uc}}$ is a monoidal category.
@@
Since $X,X^{\prime}$ are Polish, $\Sigma_{X}\otimes\Sigma_{X^{\prime}}=\mathcal{B}(X\times X^{\prime})$.
1. Tensor is a bifunctor. For morphisms $F:X\Rightarrow Y$ and $G:X^{\prime}\Rightarrow Y^{\prime}$ the map $(F\otimes G)(x,x^{\prime})=F(x)\times G(x^{\prime})$ is
@@
Weakly measurable. Let $O=\bigcup_{k\in\mathbb{N}}(B_{k}\times C_{k})$ be an open subset of $Y\times Y^{\prime}$ written as a countable union of basic rectangles (recall that $Y,Y^{\prime}$ are second-countable, so any open set in $Y\times Y^{\prime}$ is a countable union of basic rectangles). Then,
@@
2. Associativity, unit, symmetry. Because $\otimes$ is the cartesian product on objects and the point-wise cartesian product on fibers, we have the canonical homeomorphisms
@@
Each of these maps is single-valued and continuous, hence a morphism in $\mathbf{WMeas}_{\text{uc}}$ (singleton fibers are compact). They play the role of associator, left- and right-unitors, and symmetry, respectively.
3. Coherence. The associator, unitors, and symmetry just introduced are the usual homeomorphisms that witness the cartesian product as a symmetric monoidal structure in Set. Consequently the pentagon, triangle, and hexagon coherence diagrams commute exactly as they do in Top (hence also in Meas), and the same proofs apply verbatim in $\mathbf{WMeas}_{\text{uc}}$.
@@ 6.2. Functors Between Categories
Functors do exist if we focus on restrictions of $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$. [^14] We defer the study of monoidal functors to future research. In particular, call $\mathbf{WMeas}_{\text{uc,uhc}}\subset\mathbf{WMeas}_{\text{uc}}$ the category whose objects are compact Polish spaces with the Borel $\sigma$-algebra, and whose morphisms are uniformly compact-valued, weakly measurable, and upper hemicontinuous correspondences. Then, the following holds.
@@
(2) Composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be morphisms in $\mathbf{WMeas}_{\text{uc,uhc}}$.
Uniform compact bound. If $K_{\Psi}\subset Z$ is a uniform compact bound for $\Psi$, then
$(\Psi\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\Psi(y)\subset K_{\Psi}\quad\text{for all }x\in X,$
so $\Psi\circ\Phi$ is uniformly bounded.
Compact fibers. Since $\Phi(x)$ is compact, $\Psi$ has compact values, and $Z$ is compact Hausdorff, upper hemicontinuity of $\Psi$ implies $\operatorname{Gr}(\Psi)\subset Y\times Z$ is closed (by Lemma 3). Hence
$G_{x}\coloneqq\operatorname{Gr}(\Psi)\cap\bigl{(}\Phi(x)\times K_{\Psi}\bigr{)}$
is compact; Its projection onto $Z$ equals $(\Psi\circ\Phi)(x)$, so each $(\Psi\circ\Phi)(x)$ is compact.
Upper hemicontinuity. By (Aliprantis and Border,, 2006, Theorem 17.23), with $\Phi$ and $\Psi$ both u.h.c., the composition $\Psi\circ\Phi$ is u.h.c.
Weak measurability. Preserved under composition as in the proof of Theorem 23 (via Castaing selectors).
Thus $\Psi\circ\Phi$ is again a morphism in $\mathbf{WMeas}_{\text{uc,uhc}}$, and $\mathcal{F}(\Psi\circ\Phi)=\mathcal{F}(\Psi)\circ\mathcal{F}(\Phi)$.
@@
Let now $\mathbf{UHCont}_{\text{c}}^{\text{P}}\subset\mathbf{UHCont}$ be the category whose objects are compact Polish topological spaces, and whose morphisms are upper hemicontinuous, uniformly compact-valued correspondences. Then, the following holds.
Theorem 17 (Faithful Functor between $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{P}}$ and $\mathbf{WMeas}_{\text{uc}}$).
@@
$\mathcal{G}:\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{P}}\rightarrow\mathbf{WMeas}_{\text{uc}}$
@@
If $X$ is compact Polish, then $(X,\tau_{X},\mathcal{B}(\tau_{X}))$ is a compact Polish measurable space, so $\mathcal{G}$ is well-defined on objects.
Let $\Phi:X\rightrightarrows Y$ be a morphism in $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{P}}$, i.e., upper hemicontinuous with uniformly compact values (hence closed values). Since $Y$ is compact Polish, $\Phi$ is automatically uniformly compact-valued (take $K_{\Phi}=Y$). We now show $\Phi$ is weakly measurable.
Weak measurability. Fix open $O\subseteq Y$ and define $\varphi_{O}(y)\coloneqq d(y,O^{c})$, where $d$ is the metric associated with the topology on $Y$. Set $h_{O}(x)\coloneqq\sup_{y\in\Phi(x)}\varphi_{O}(y)$. By Berge’s maximum theorem $h_{O}$ is upper semicontinuous. Hence
$\{x:\Phi(x)\cap O\neq\emptyset\}=\{x:h_{O}(x)>0\}=\bigcup_{m=1}^{\infty}\{x:h_{O}(x)\geq 1/m\},$
a countable union of closed sets; therefore it is measurable. Hence $\Phi$ is weakly measurable. Therefore $\mathcal{G}(\Phi)$ is a morphism in $\mathbf{WMeas}_{\mathrm{uc}}$.
Functoriality. Identities are preserved trivially. For composition, if $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ are upper hemicontinuous, then $\Psi\circ\Phi$ is upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23). Because $\Psi$ is upper hemicontinuous with compact values and $Z$ is compact Hausdorff, Lemma 3 implies that $\operatorname{Gr}(\Psi)\subset Y\times Z$ is closed. For each $x\in X$, set $G_{x}\coloneqq\operatorname{Gr}(\Psi)\cap\bigl{(}\Phi(x)\times Z\bigr{)}.$ Since $\Phi(x)$ is compact and $Z$ is compact Hausdorff, the product $\Phi(x)\times Z$ is compact; Hence $G_{x}$ is compact as a closed subset of a compact set. The projection onto $Z$ satisfies
$\pi_{Z}(G_{x})=\bigcup_{y\in\Phi(x)}\Psi(y)=(\Psi\circ\Phi)(x),$
and since $\pi_{Z}:Y\times Z\to Z$ is continuous, $\pi_{Z}(G_{x})$ is compact. Therefore $(\Psi\circ\Phi)(x)$ is compact, and the same uniform bound $K_{\Psi}=Z$ works. Weak measurability of $\Psi\circ\Phi$ follows from the previous argument applied to the composite. Thus $\mathcal{G}(\Psi\circ\Phi)=\mathcal{G}(\Psi)\circ\mathcal{G}(\Phi)$.
Finally, $\mathcal{G}$ is the identity on morphisms, so it is injective on each $\mathrm{Hom}$-set and hence faithful. ∎
@@ 6.3. Monads
Theorem 19 (Vietoris lax monad on $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$).
Define an endofunctor $T$ and morphisms $\eta,\nu$ as follows:
  • On objects: $T(X)\coloneqq\mathcal{K}(X)$, the space of nonempty compact subsets of $X$ with the Vietoris topology.
  • On morphisms: for $\Phi:X\rightrightarrows Y$ and $K\in\mathcal{K}(X)$, set
    $T(\Phi)(K)\coloneqq\bigl{\{}L\in\mathcal{K}(Y):L\subset\Phi[K]\bigr{\}},\qquad\Phi[K]\coloneqq\bigcup_{x\in K}\Phi(x).$
  • Unit: $\eta_{X}:X\to\mathcal{K}(X)$, $\eta_{X}(x)=\{x\}$ (single-valued morphism).
  • Multiplication: $\nu_{X}:\mathcal{K}(\mathcal{K}(X))\rightrightarrows\mathcal{K}(X)$,
    $\nu_{X}(\mathcal{A})=\Bigl{\{}\bigcup_{A\in\mathcal{A}^{\prime}}A:\emptyset\neq\mathcal{A}^{\prime}\subseteq\mathcal{A}\text{ compact in }\mathcal{K}(X)\Bigr{\}},$
where $\mathcal{K}(\mathcal{K}(X))$ denotes the hyperspace whose elements are nonempty compact subsets of $\mathcal{K}(X)$. Then, $T$ is an endofunctor on $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$, $\eta$ and $\nu$ are natural (single-valued, resp. set-valued) morphisms, and $(T,\eta,\nu)$ satisfies the lax monad laws (with respect to inclusion of graphs),
$\mathrm{id}\subseteq\nu\circ T\eta,\qquad\nu\circ\eta T=\mathrm{id},\qquad\nu\circ T\nu\subseteq\nu\circ\nu T.$
Moreover, on the subcategory of single-valued continuous maps (CHaus), these inclusions become equalities, recovering the classical Vietoris monad.
@@
1) $T$ is a well-defined endofunctor. For a morphism $\Phi:X\rightrightarrows Y$, values of $T(\Phi)$ are nonempty compact subsets of $\mathcal{K}(Y)$. Indeed, $K\mapsto\Phi[K]$ is compact for each compact $K\subset X$ (since $\Phi$ has a closed graph and compact values, and $K$ is compact), so $T(\Phi)(K)=\mathcal{K}(\Phi[K])$ is compact in $\mathcal{K}(Y)$. To see upper hemicontinuity, it suffices to show the graph of $T(\Phi)$ is closed in $\mathcal{K}(X)\times\mathcal{K}(Y)$. Let $K_{m}\to K$ in $\mathcal{K}(X)$ and $L_{m}\to L$ in $\mathcal{K}(Y)$ with $L_{m}\subset\Phi[K_{m}]$. Pick any $y\in L$; Choose $y_{m}\in L_{m}$ with $y_{m}\to y$. For each $m$, pick $x_{m}\in K_{m}$ with $y_{m}\in\Phi(x_{m})$. By compactness of $X$ and the Vietoris convergence $K_{m}\to K$, the net $(x_{m})$ has a subnet $(x_{m_{\ell}})$ converging to some $x\in K$. Upper hemicontinuity of $\Phi$ then gives $y\in\Phi(x)$, hence $y\in\Phi[K]$. As $y\in L$ was arbitrary, $L\subset\Phi[K]$; Thus $(K,L)\in\operatorname{Gr}(T(\Phi))$. Therefore $T(\Phi)$ is compact-valued, upper hemicontinuous. Functoriality is immediate,
$T(\Psi\circ\Phi)(K)=\{L:L\subset(\Psi\circ\Phi)[K]\}$ $=\{L:L\subset\Psi[\Phi[K]]\}$
$=\bigcup_{M\in\mathcal{K}(\Phi[K])}\{L\in\mathcal{K}(Y):L\subset\Psi[M]\}=(T\Psi\circ T\Phi)(K),$
and $T(\mathrm{id})=\mathrm{id}$.
2) $\eta$ and $\nu$ are morphisms (natural in $X$). For $\eta_{X}$, singletons are compact and the map $x\mapsto\{x\}$ is continuous into $\mathcal{K}(X)$, hence a morphism. For $\nu_{X}$, values are compact: If $\mathcal{A}^{\prime}\subset\mathcal{A}$ is compact in $\mathcal{K}(X)$, then $\bigcup_{A\in\mathcal{A}^{\prime}}A$ is a compact subset of $X$. Upper hemicontinuity of $\nu_{X}$ follows from the Vietoris calculus: For a basic open set $\langle U_{1},\dots,U_{m}\rangle\subset\mathcal{K}(X)$,
$\nu_{X}^{-1}\bigl{(}\langle U_{1},\dots,U_{m}\rangle\bigr{)}\supseteq\Bigl{\langle}[U_{1}],\dots,[U_{m}],\langle U_{1}\cup\cdots\cup U_{m}\rangle\Bigr{\rangle},$
and the right hand side is open in $\mathcal{K}(\mathcal{K}(X))$ (see Definition 28); Hence $\nu_{X}$ is upper hemicontinuous.
Naturality (in the lax sense) holds by direct computation: For any $\Phi:X\rightrightarrows Y$,
$(T\Phi)\circ\eta_{X}\supseteq\eta_{Y}\circ\Phi\qquad\text{and}\qquad\nu_{Y}\circ T(T\Phi)\supseteq T(\Phi)\circ\nu_{X},$
since $\{y\}\subset\Phi(x)$ iff $y\in\Phi(x)$, and unions commute with direct images.
3) Lax monad laws. All inclusions are understood as inclusions of correspondences (i.e., of their graphs).
Left unit (lax): For $K\in\mathcal{K}(X)$,
$(T\eta_{X})(K)=\bigl{\{}\mathcal{C}\in\mathcal{K}(\mathcal{K}(X)):\mathcal{C}\subset\eta_{X}[K]\bigr{\}}.$
Thus